An M30 × 3.5 ISO grade 8.8 bolt is used in a joint at a preload level recommended for reused connections, and the joint is subject to a repeated tensile fatigue load of P = 65 kN per bolt. The joint constant is C = 0.25.

  1. Determine the static load factors;                                                                   [12 pts.]
  2. The factor of safety guarding against a fatigue failure based on the Goodman fatigue criterion.                                                                                                     [18 pts.]
Answer #1

M30 * 35: from standard tables tensile stress area (Ax ) = 5 6L mm2 for M16-36 strength (Sut 830 mla Minie proof strength (8)@ the yielding factor of safety Goox 561 no = Sp. At cp+fi (0:25*6500) +859450 mo 1095 load factor n = SpAt fi CP M = (600x56factor of safety for Goodman fatigue criterion 6 n Se ( Sut-05) ca (Sut +Se Me = 199 * (830-450) 14.48*(830 +129) Mf = 3.53